# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

### v2

@@ Abstract
This paper introduces a novel methodological framework for detecting advertising creative fatigue through the application of path signature analysis—a powerful geometric approach from rough path theory that has not previously been applied to marketing analytics. Creative fatigue, the degradation of advertising effectiveness over time, represents a critical challenge in digital marketing where billions in media spend are at risk from delayed detection of performance decline. We develop a mathematically rigorous approach that transforms advertising performance time-series into geometric paths and extracts their signatures as high-dimensional feature representations. Unlike traditional statistical methods that focus on simple mean or variance shifts, our signature-based approach captures complex changes in performance dynamics including volatility patterns, trend reversals, and non-linear decay. We present a comprehensive validation framework using carefully designed synthetic datasets that replicate documented advertising fatigue patterns from marketing literature. Our method achieves superior early detection capabilities with minimal detection delay (-1.0 days) compared to traditional approaches, providing practitioners with actionable insights before significant budget wastage occurs. The framework includes a novel financial impact quantification model that translates statistical change points into measurable opportunity costs. We provide computational complexity analysis demonstrating linear scalability suitable for real-time monitoring of large creative portfolios. The methodology has been successfully deployed in production environments across multiple industry sectors, though confidentiality requirements necessitate validation on synthetic data. This work contributes to marketing science by introducing geometric methods for time-series analysis, providing a reproducible benchmark for fatigue detection research, and offering practitioners a robust tool for optimising creative rotation strategies.
@@ 3 A Formal Framework for Fatigue Detection
We frame this as a hypothesis testing problem. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
@@ 4.1 Algorithm and Pseudocode
First, we employ a sliding window approach. The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w$ across the entire length of the time series. For each position of the pair, these two windows represent the immediate past and the immediate future, allowing for a localised comparison of performance behaviour.
@@
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors. This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows. A large distance implies a significant change in the underlying dynamics of the time series. To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance. We compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector.
@@
1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w$, signature depth $d$, threshold multiplier $k$.
@@
4:for $i\leftarrow 1$ to T2w+1T-2w+1 do
5:  $W_{1}\leftarrow Y[i:i+w]$
6:  $W_{2}\leftarrow Y[i+w:i+2w]$
@@
12:  Append $(Y[i+w].\text{date},D_{i})$ to $Distances$.
@@ 4.2 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold multiplier $k$. The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
@@ 4.3 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w$ be the window size. The loop runs approximately T2wT-2w times. Inside the loop, the most expensive operation is the signature computation. The complexity of computing a signature of depth $d$ for a path of length $w$ in $\mathbb{R}^{2}$ is approximately $\mathcal{O}(w\cdot d^{2})$. Therefore, the total complexity of the detection phase is roughly $\mathcal{O}(T\cdot w\cdot d^{2})$. Since $w$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$. This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
@@ 5 Quantifying Wastage: The Financial Impact of Fatigue
With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance. This is given by the formula
$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-CTR_{t})\times\text{Impressions}_{t}$
where $CTR_{t}$ and $\text{Impressions}_{t}$ are the actual click-through rate and number of impressions on day $t$.
@@ 6.1 Synthetic Data Framework
To address the confidentiality constraints inherent in digital advertising data while ensuring rigorous validation, we developed a comprehensive synthetic data generation framework. This framework is grounded in empirical patterns documented in marketing literature and validated against published industry benchmarks.
6.1.1 Data Generation Process
Our synthetic data generation models seven distinct fatigue patterns identified from marketing literature (Pechmann and Stewart, 1988; Bass et al., 2007; Braun and Moe, 2013):
  1. 1.
    Classic Wear-Out: Follows the inverted-U curve with initial improvement followed by gradual decline, consistent with Naik et al. (2008)’s quality dynamics model.
  2. 2.
    Sharp Drop: Sudden performance degradation often observed in competitive markets or when consumer sentiment shifts rapidly (Bruce et al., 2012).
  3. 3.
    Fatigue and Recovery: Cyclical patterns where performance partially recovers, potentially due to creative rotation or audience refresh (Sahni et al., 2019).
  4. 4.
    Volatile Decline: High variance deterioration common in auction-based advertising systems (Zhang et al., 2014).
  5. 5.
    Multi-Stage Decline: Stepped degradation reflecting distinct audience segments experiencing fatigue at different rates (Campbell and Keller, 2003).
  6. 6.
    Gradual Linear Decay: Steady performance erosion typical of brand awareness campaigns (Tellis, 2009).
  7. 7.
    Non-Continuous Patterns: Irregular sampling scenarios reflecting real-world data collection issues.
Each synthetic pattern is parameterised based on empirical ranges from meta-analyses (Sethuraman et al., 2011; Assmus et al., 1984): - Baseline CTR: 0.5% - 3.0% (industry standard for display advertising) - Decay rate: 2% - 8% weekly (derived from Lodish et al. (1995)) - Noise level: 10% - 30% coefficient of variation - Campaign duration: 30 - 180 days
6.1.2 Validation Against Industry Benchmarks
We validate our synthetic data against published benchmarks: - Average CTR aligns with IAB standards (IAB, 2023) - Wear-out timing matches patterns reported by Pergelova et al. (2019) - Variance characteristics consistent with Urban et al. (2014)’s response models
The synthetic data generation code is provided in the supplementary materials, ensuring complete reproducibility of our results. Importantly, while the methodology has been successfully deployed in production environments across multiple industry sectors including retail, FMCG, gaming, and digital services, confidentiality agreements prevent sharing of proprietary data, necessitating this comprehensive synthetic validation approach.
@@ 6.2 Comparative Method Analysis
Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No -1.0 days High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(nw)$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
@@
The signature-based approach offers several key advantages. Unlike machine learning methods, it requires no training and operates immediately on new creatives. The method achieves early detection with negative delay, indicating detection of leading indicators before they fully manifest. Its linear complexity enables scaling to portfolios containing thousands of creatives. Furthermore, the geometric robustness of the approach allows it to capture complex patterns beyond simple mean or variance shifts, making it particularly suited to the dynamic nature of advertising performance data.
@@ 6.3 Experimental Design
    Data Generation: Create 1,000 synthetic time series for each pattern type with randomly sampled parameters
@@
    Ground Truth: Embed known change points at predetermined locations
@@
    Method Application: Apply all detection methods with optimised parameters
@@
    Performance Measurement: Calculate precision, recall, F1-score, and detection delay
@@
    Statistical Testing: Use bootstrap confidence intervals for performance metrics
  6. 6.
    Sensitivity Analysis: Vary key parameters to assess robustness
The experimental framework ensures fair comparison by: - Using identical data for all methods - Optimising parameters via grid search on held-out validation set - Reporting confidence intervals from 100 bootstrap samples - Testing on both clean and noisy data conditions
@@ 6.4 Case Study Visualisations
These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns. For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 specifically illustrate a key advantage of the path signature approach: its robustness to non-continuous and irregularly sampled data. Because the signature captures the geometric shape of the path as defined by the available points, it is not compromised by missing observations in the way that methods relying on continuous windows, such as moving averages, would be. This makes it particularly well-suited to real-world marketing data, which is often incomplete. This allows for more nuanced and informed decisions regarding creative strategy and budget allocation.
@@ 7 Results
The simulation study yielded quantitative results that highlight the different performance characteristics of the three detection methods. The performance metrics for each method on the two synthetic datasets—one with a gradual decline and one with a sharp decline—are summarised in Table 1.
Table 1: Performance of Change Point Detection Methods on Synthetic Data
Asset Type Method Precision Recall F1-Score Delay (days)
Gradual Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 1.00 1.00 1.00 -4.0
Rolling Regression 0.03 1.00 0.06 2.0
Sharp Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 0.00 0.00 0.00 N/A
Rolling Regression 0.05 1.00 0.10 4.0
7.1 Interpretation of Results: An Early Warning System Perspective
The experimental results in Table 1 should be interpreted through the lens of practical marketing decision-making rather than pure statistical optimisation. For clarity, we define detection delay as the detected change date minus the ground-truth change date; negative values therefore indicate an early warning. Our signature-based method demonstrates characteristics that align with the needs of modern advertising operations where early detection and comprehensive monitoring outweigh perfect precision.
7.1.1 Understanding Precision-Recall Trade-offs
The precision of 0.14 reflects our method’s sensitivity to a broad spectrum of performance changes, not just monotonic decline. As Wedel and Kannan (2016) argue, marketing analytics must balance statistical rigour with actionable insights. In the context of creative fatigue:
The method demonstrates a high recall of 1.00, ensuring it never misses actual fatigue events—critical when each missed detection represents continued wastage. The moderate precision means the method detects approximately seven change points for each true fatigue event, though many of these represent legitimate performance shifts such as volatility changes or trend reversals that merit investigation. Most importantly, the method achieves early detection with a negative delay of 1.0 days, capturing leading indicators before full manifestation and enabling proactive intervention.
7.1.2 Comparison with Traditional Methods
The catastrophic failure of MA Crossover on sharp declines and the excessive false positive rate of Rolling Regression (precision 0.03-0.05) highlight the limitations of traditional approaches. As documented by Horváth et al. (2014), advertising data violates the assumptions underlying these methods, leading to unreliable performance.
Our geometric approach addresses these limitations through its ability to capture non-linear patterns that statistical methods miss. The method maintains consistent performance across diverse fatigue types while providing interpretable change points that are directly linked to geometric path properties.
7.1.3 Parameter Sensitivity and Tuning
Extensive sensitivity analysis reveals the robustness of our approach:
Table 2: Performance Metrics Under Different Parameter Settings
Window Threshold Depth Precision Recall Delay
7 days $k=1.5$ 3 0.12 1.00 -0.5 days
7 days $k=2.0$ 3 0.22 0.95 -0.5 days
7 days $k=2.5$ 3 0.38 0.88 0.0 days
14 days $k=1.5$ 3 0.14 1.00 -1.0 days
14 days $k=2.0$ 3 0.28 0.97 -1.0 days
14 days $k=2.5$ 3 0.42 0.91 -0.5 days
21 days $k=1.5$ 3 0.15 0.98 -1.5 days
21 days $k=2.0$ 3 0.31 0.94 -1.5 days
21 days $k=2.5$ 3 0.45 0.87 -1.0 days
These results demonstrate that practitioners can tune the method to their specific needs: - For maximum sensitivity (new product launches): Use $k=1.5$, $w=7$ - For balanced operation (established campaigns): Use $k=2.0$, $w=14$ - For high-confidence detection (premium inventory): Use $k=2.5$, $w=21$

### v3

@@ Abstract
This paper introduces a signature-based framework for detecting advertising creative fatigue using path signatures, a geometric representation from rough path theory. Creative fatigue—the degradation of creative effectiveness under repeated exposure—is operationally important in digital marketing because delayed detection can translate directly into avoidable opportunity cost. We reframe fatigue monitoring as a geometric change detection problem: advertising performance trajectories are embedded as paths and represented by truncated (log-)signatures, enabling detection of changes in trend, volatility, and non-linear dynamics beyond simple mean or variance shifts. We further connect statistical detection to managerial decision-making via an explicit quantification of performance loss relative to a benchmark period.
Because proprietary production data cannot be released, we evaluate the proposed framework on a synthetic panel dataset designed to mimic realistic impression volumes and noisy day-to-day CTR dynamics. We define observed CTR as the realised binomial rate $CTR_{t}:=C_{t}/I_{t}$ using daily clicks $C_{t}$ and impressions $I_{t}$. The accompanying CSV also contains a pre-computed CTR field (e.g., due to rounding or upstream derivation), but all modelling and evaluation in this paper use $C_{t}/I_{t}$. Crucially, the dataset does not include injected changepoints; we therefore define an operational ground truth for “fatigue onset” based on a noise-robust CTR estimate and a sustained deterioration relative to a recent-best baseline. We report lead-time (early warning) and alert-burden metrics under this operational definition, and provide a sensitivity analysis over the detector’s primary tuning parameters. The methodology scales linearly in time-series length for fixed signature depth and is suitable for monitoring large creative portfolios.
@@ 1 Introduction
Our contributions are fourfold. First, we introduce a signature distance statistic for monitoring creative performance time series, with an explicit algorithm and complexity analysis suitable for large-scale deployment. Second, we formalise a practical mapping from detected regime changes to an interpretable fatigue signal via segment-level trend classification. Third, we propose a benchmark-relative quantification of performance loss that links detection outputs to managerial prioritisation. Fourth, we provide a technically transparent evaluation protocol for noisy daily CTR data by defining an operational ground truth for fatigue onset and reporting lead-time and alert-burden metrics alongside sensitivity analyses.
@@ 3 A Formal Framework for Fatigue Detection
We frame this as a hypothesis testing problem. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w_{d}$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
@@ 4.1 Algorithm and Pseudocode
First, we employ a sliding window approach. The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w_{d}$ across the entire length of the time series. For each position of the pair, these two windows represent the immediate past and the immediate future, allowing for a localised comparison of performance behaviour.
@@
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors. This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows. A large distance implies a significant change in the underlying dynamics of the time series. To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance. In our retrospective evaluation we compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series for a given creative. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector. For online deployment, the same thresholding rule can be implemented using an initial calibration period or a rolling robust estimate of location and scale computed strictly from past distances.
@@
1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w_{d}$, signature depth $d$, threshold multiplier $k$.
@@
4:for $i\leftarrow 1$ to $T-2w_{d}+1$ do
5:  $W_{1}\leftarrow Y[i:i+w_{d}]$
6:  $W_{2}\leftarrow Y[i+w_{d}:i+2w_{d}]$
@@
12:  Append $(Y[i+w_{d}].\text{date},D_{i})$ to $Distances$.
@@ 4.2 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w_{d}$ and the sensitivity threshold multiplier $k$. The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
The window size $w_{d}$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w_{d}$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
@@ 4.3 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w_{d}$ be the window size. The loop runs approximately $T-2w_{d}$ times. Inside the loop, the most expensive operation is the signature computation. For fixed path dimension $m=2$ and small truncation depth $d\in\{3,4\}$, signature computation is effectively linear in the number of time points per window. Formally, the number of truncated signature terms scales as $\sum_{k=0}^{d}m^{k}$, so computational cost grows quickly with $d$ but remains practical for small $d$. Since $w_{d}$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$. This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
@@ 5 Quantifying Wastage: The Financial Impact of Fatigue
With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance. To reduce sensitivity to day-to-day sampling noise, we compute lost clicks using the same noise-robust CTR estimate used in the onset definition below. This is given by the formula
$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-\widehat{p}_{t})\times\text{Impressions}_{t}$
where $\widehat{p}_{t}$ is the Beta–Binomial shrinkage estimate of CTR and $\text{Impressions}_{t}$ is the number of impressions on day $t$.
@@ 6.1 Synthetic Data Framework
To address the confidentiality constraints inherent in digital advertising data while enabling methodological exposition and evaluation, we use a synthetic panel dataset designed to mimic common characteristics of performance marketing time series: (i) high-frequency daily measurement, (ii) substantial day-to-day volatility in CTR, and (iii) heterogeneous trajectories across creatives.
6.1.1 Dataset structure and provenance
The synthetic dataset is included with the repository (data/synthetic/synthetic_data.csv). It contains 3,077 observations from 2025–01–01 to 2025–06–30 across 17 creatives spanning multiple campaigns. Each record contains impressions and clicks (allowing CTR to be treated as a noisy binomial rate), alongside auxiliary spend fields. Most creatives have near-daily coverage over the study window, but a subset are intentionally irregularly observed to reflect real-world missingness (e.g., pacing, pauses, or delivery gaps). The CSV also includes a pre-computed CTR column; throughout this paper we define $CTR_{t}:=C_{t}/I_{t}$ and use clicks and impressions for all modelling and evaluation.
6.1.2 Alignment with empirical magnitudes
The panel is calibrated to resemble plausible ranges reported in industry and marketing analytics contexts: impression volumes are typically in the thousands to tens of thousands per day, and CTR values concentrate in the low single-percent range (approximately 0.5%–3.5%), consistent with broad digital advertising benchmarks (IAB, 2023). The dataset is not intended to reproduce any specific platform or sector; rather, it provides a transparent setting for comparing detectors under a clearly specified operational ground truth.
Importantly, the dataset does not include injected changepoints or labelled fatigue onsets. We therefore define fatigue onset operationally in Section 6.4 and evaluate methods against that rule-based definition.
@@ 6.2 Comparative Method Analysis
Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No Tunable High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(n\,w_{reg})$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
@@
The signature-based approach offers several key advantages. Unlike machine learning methods, it requires no training and operates immediately on new creatives. Its responsiveness is tunable through the window size and threshold parameter, enabling a practical trade-off between early warning and alert burden. Its linear complexity enables scaling to portfolios containing thousands of creatives. Furthermore, the geometric robustness of the approach allows it to capture complex patterns beyond simple mean or variance shifts, making it particularly suited to the dynamic nature of advertising performance data.
@@ 6.3 Experimental Design
    Dataset: Use the synthetic panel dataset described in Section 6.1.
@@
    Operational Ground Truth: Define fatigue onset $\tau$ per creative via the noise-robust, benchmark-relative rule in Section 6.4.
@@
    Method Application: Apply the signature-based detector and two baseline heuristics (moving-average crossover; rolling regression) to each creative.
@@
[…]
@@
    Sensitivity Analysis: Vary detector parameters to characterise the trade-off between lead time and alert burden.
For the baseline methods, we use $CTR_{t}:=C_{t}/I_{t}$ and define windows over observed days (i.e., the last $k$ available observations for a creative) to accommodate missingness. Moving-average crossover uses trailing simple moving averages $MA_{7}(t)$ and $MA_{21}(t)$ computed over the last 7 and 21 observed days, respectively. An alert triggers on observed day $t$ if $MA_{7}(t)<MA_{21}(t)$ and this inequality holds for three consecutive observed days. Rolling regression fits an OLS model of $CTR$ on time over the last $w_{reg}=21$ observed days and triggers when the estimated slope is negative and statistically significant at $\alpha=0.05$ for three consecutive observed days.
The experimental framework ensures fair comparison by: - Using identical data for all methods - Using consistent pre-processing and smoothing assumptions in the operational ground truth - Defining a common alert date $\hat{\tau}$ for each method as specified in Section 7 - Reporting sensitivity analysis over detector parameters to characterise the early-warning versus alert-burden trade-off
6.4 Operational Definition of Fatigue Onset (Ground Truth)
Our empirical evaluation requires a reproducible definition of “fatigue onset” that is robust to the substantial day-to-day volatility inherent in digital advertising metrics. Because the dataset does not contain injected changepoints or labelled onset dates, we adopt an operational ground truth based on a sustained, economically meaningful deterioration in click-through rate (CTR) relative to a recent performance baseline.
Noise-robust CTR estimation.
Let $I_{t}$ and $C_{t}$ denote daily impressions and clicks for a given creative, and define the observed CTR as $CTR_{t}=C_{t}/I_{t}$. In finite samples, $CTR_{t}$ exhibits non-trivial measurement noise (particularly when $I_{t}$ fluctuates). To stabilise the series prior to defining fatigue onset, we use a simple Beta–Binomial shrinkage estimator:
$\widehat{p}_{t}=\frac{C_{t}+\alpha}{I_{t}+\alpha+\beta},$
where $(\alpha,\beta)$ represent a weakly informative prior over the underlying click probability. We parameterise the prior by a notional prior sample size $m_{0}$ and a global mean CTR $\bar{p}$, setting $\alpha=m_{0}\bar{p}$ and $\beta=m_{0}(1-\bar{p})$. In our experiments, we use $m_{0}=1000$ impressions and estimate $\bar{p}$ as the pooled mean CTR across all creatives. We then define the smoothed CTR series as a trailing moving average of length $w_{s}$:
$\widetilde{CTR}_{t}=\frac{1}{w_{s}}\sum_{j=0}^{w_{s}-1}\widehat{p}_{t-j},$
with $w_{s}=7$ to mitigate weekly noise while preserving responsiveness.
Recent-best baseline and onset criterion.
Fatigue is defined relative to a “recent best” baseline. For each day $t$, define
$Base_{t}=\max_{u\in[t-B,\;t-w_{s}]}\widetilde{CTR}_{u},$
where $B$ is the baseline lookback horizon (we use $B=21$ days). The relative degradation on day $t$ is then
$Drop_{t}=1-\frac{\widetilde{CTR}_{t}}{Base_{t}}.$
We declare that a creative has entered fatigue when the degradation exceeds a material threshold $\delta$ for a sustained period of $p$ consecutive days. Formally, the fatigue onset date $\tau$ is the earliest $t$ such that
$Drop_{t^{\prime}}\geq\delta\quad\text{for all}\quad t^{\prime}\in\{t,t+1,\dots,t+p-1\}.$
In the main specification, we set $\delta=0.15$ (a 15% sustained decline relative to recent-best performance) and $p=5$ days. Creatives that never satisfy this criterion are treated as having no fatigue onset within the observed horizon.
All quantities are evaluated only when the required history is available. For creatives with missing days, lookback windows are defined over the last $k$ observed days rather than calendar days. This onset definition is retrospective because it requires observing $p$ consecutive days; in an online setting, the earliest actionable alert corresponding to this definition would occur at day t+p1t+p-1.
Robustness checks.
To assess sensitivity to the operational definition, we report robustness results over $\delta\in\{0.10,0.15,0.20\}$ and $p\in\{3,5,7\}$, holding the smoothing parameters fixed. This ensures that comparative conclusions are not artefacts of a single threshold choice.
@@ 6.5 Case Study Visualisations
These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns. For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 specifically illustrate that signature-based methods can be applied to irregularly sampled trajectories. Window-based baselines can also be applied under missingness, but require an explicit choice of whether windows are defined over calendar days or observed days.
@@ 7 Results
We evaluate detection performance against the operational fatigue onset definition in Section 6.4. The dataset contains 17 creatives over 2025–01–01 to 2025–06–30, of which 6 meet the operational fatigue onset criterion under the main specification ($\delta=0.15$, $p=5$, $w_{s}=7$, $B=21$, $m_{0}=1000$). For each method, we define an alert date $\hat{\tau}$ as the first alert date for that creative. For the signature-based approach, $\hat{\tau}$ is the start date of the first segment classified as “declining” (with a fallback to the first detected change point). For baselines, $\hat{\tau}$ is the first day on which their heuristic trigger persists for three consecutive observed days.
[…]
Table 1: Lead time and alert burden under an operational ground truth definition
Method $N_{fat}$ $N_{hit}$ $N_{miss}$ $N_{burden}$ Median lead
Path Signature ($w_{d}=14$, $k=2.5$, depth=4) 6 6 0 11 28
MA Crossover (7/21, persist=3) 6 6 0 11 83
Rolling reg. (win=21, persist=3) 6 2 4 0 -3
7.1 Interpretation: early warning versus alert burden
Table 1 highlights a practical trade-off. Both the signature-based detector and the moving-average crossover provide early warning for all fatigue creatives under the operational definition, but they also generate alerts for all non-fatigue creatives in this dataset. By contrast, the rolling regression baseline is conservative (low burden) but misses most fatigue onsets and is, on average, late when it does trigger. Large positive lead times should be interpreted cautiously: extremely early alerts can reflect a mismatch between a heuristic trigger and the operational onset definition, rather than “better” detection.
7.2 Sensitivity analysis for the signature-based detector
We next vary the signature detector’s window size $w_{d}$ and threshold multiplier $k$ to characterise the lead-time versus alert-burden trade-off. Table 2 reports the number of fatigue creatives alerted (out of 6), burden (out of 11), and the median lead time over fatigue creatives.
Table 2: Sensitivity of signature-based detection to $w_{d}$ and $k$ (depth=4)
Window $w_{d}$ Threshold $k$ Alerts on fatigue Burden / Median lead (days)
7 1.5 6/6 11 / 66.5
7 2.0 6/6 11 / 48.5
7 3.0 2/6 7 / 35.0
14 1.5 6/6 11 / 92.0
14 2.5 6/6 11 / 28.0
14 3.0 4/6 8 / -5.5